OpenStudy (miss.rose):

@Ghostgate @JoeDeWise @rebeccaxhawaii Can you help me with five questions on "The Distributive Property" I will medal! Please help!

rebeccaxhawaii (rebeccaxhawaii):

yes post

OpenStudy (benlindquist):

Which are the questions

OpenStudy (anonymous):

example: 2*(4+5+6) = 2*4 + 2 * 5 + 2 * 6.

rebeccaxhawaii (rebeccaxhawaii):

okay so first multiply 2 by 4, 5, and 6

OpenStudy (miss.rose):

The first question is: Which expression is equivalent to this one? –5(6 + 11) A. (–5) · 6 + 11 B. (–5) + 6 + 11 C. (–5) · 6 · 11 D. (–5) · 6 + (–5) · 11 This one?

rebeccaxhawaii (rebeccaxhawaii):

so 8+10+12= 8+10+ 12

OpenStudy (benlindquist):

? = ? is that all

rebeccaxhawaii (rebeccaxhawaii):

yes D is correct

OpenStudy (miss.rose):

Next Question: What is the solution to the equation? (3 + 7) = 4 · x + 4 · 7 A. x = 3 This one? B. x = 4 C. x = 7 D. x = 10

rebeccaxhawaii (rebeccaxhawaii):

okay well lets parenthesis first so 10= 4 times x +4 times 7

OpenStudy (miss.rose):

So it's ten?

rebeccaxhawaii (rebeccaxhawaii):

no because x isnt isolated

OpenStudy (miss.rose):

Okay. Now I am getting confused. What would it be?

rebeccaxhawaii (rebeccaxhawaii):

okay look so P.E.M.D.A.S parenthesis then exponents then multiplication and/or division the addition and/or subtation

rebeccaxhawaii (rebeccaxhawaii):

so thats why we got 10

rebeccaxhawaii (rebeccaxhawaii):

\[10=4 \times x + 4 \times 7\]

rebeccaxhawaii (rebeccaxhawaii):

then multiply is next up

rebeccaxhawaii (rebeccaxhawaii):

so 10= 4x+28

rebeccaxhawaii (rebeccaxhawaii):

you are trying to get x by itself

rebeccaxhawaii (rebeccaxhawaii):

so move everything else on the other side of the equal sign

rebeccaxhawaii (rebeccaxhawaii):

so 10-28=4x the 28 was moved to the other side therefore you need to change the sign so it was positive now its negative

OpenStudy (miss.rose):

@Ghostgate @JoeDeWise What do you think?

OpenStudy (anonymous):

I think that right.

OpenStudy (anonymous):

is

rebeccaxhawaii (rebeccaxhawaii):

are u saying i dont know how to do it

OpenStudy (miss.rose):

No, not at all. I was just saying that I was trying to get another opinion because I am not understanding what you are trying to say. Nothing personal.

OpenStudy (ghostgate):

Problem 1 is D, like rebeccaxhawaii claimed, however: Problem 2: When I do the Math and distribute I come out with -4.5, which allows 10 and 10 to equal, but... the current problem isn't like that.

OpenStudy (miss.rose):

Please keep this in mind I am in the 7th grade so I have not learned how you guys are doing it.

OpenStudy (ghostgate):

Basically rebeccaxhawaii did the Math, however worded it in a more confusing manner.

rebeccaxhawaii (rebeccaxhawaii):

well thats how i learned it so

rebeccaxhawaii (rebeccaxhawaii):

thats how ill explain it

OpenStudy (miss.rose):

@rebeccaxhawaii I understand that. Everybody learns a different way. No hard feelings?

OpenStudy (ghostgate):

Anyhow, based off of the problem I don't know where the fours come from on the right, I understand where the '7' comes from, my guess is that '3' is the answer, however originally if you were to solve it algebraically(I believe) then your answer would be -4.5.

rebeccaxhawaii (rebeccaxhawaii):

lol stop trying to make it sound more complicated

OpenStudy (ghostgate):

Anyhow, like I stated I believe 3 is the answer.

OpenStudy (miss.rose):

Okay, thanks! I will submit it when I am done and let you know what happens. Next Question: Rewrite the expression by using grouping symbols. 1/2 * 4 + 1/2 *z A. 4(1/2 + z) B. z(1/2 * 4) C. 1/2 * (4 + 1/2) * z Is it this one? D. 1/2(4 + z)

rebeccaxhawaii (rebeccaxhawaii):

c would be correct

OpenStudy (ghostgate):

Yeah I agree. C. is the answer.

OpenStudy (miss.rose):

Ok. Thanks! 4th Question: Which expression is equivalent to this one? 1/6(x - 18) A. 1/6x - 18 I am guessing this one? B. 1/6x * (-18) C. 1/6x - 3 D. 1/6 + x - 18 5th Question: Evaluate this expression without using a calculator: 8 * 63 A. 489 B. 494 C. 504 This one. D. 683

rebeccaxhawaii (rebeccaxhawaii):

yes both are correct

OpenStudy (miss.rose):

Okay. Thanks! I am going to submit it and tell what I get. Then I am going to medal. Then close the question. So Bye! THANKS FOR ALL YOUR HELP!!!!!

OpenStudy (miss.rose):

I got a 3/5. That isn't so bad. But again thanks! Bye!